# Bilevel subsidy-enabled mobility hub network design with perturbed utility coalitional choice-based assignment

### v1

@@ 1 Introduction
Emergent cyberphysical mobility hubs (MHs) are a viable response to these challenges. MHs are defined as ”multimodal transport nodes that facilitate intermodal transfers by providing different mobility options in close proximity” (Miramontes et al.,, 2017). While ”transit hubs” have existed for decades as a means to integrate land use and public transport as a type of ”transit-oriented development” (see Weustenenk and Mingardo, (2023)), they differ from MHs in the literature in the last decade, which feature higher levels of both physical and digital integration of mobility services (Arias-Molinares et al.,, 2023), particularly services that feature real-time information. These MHs often involve multiple operators (or even non-mobility-related goods and service providers, such as fueling), necessitating distinct incentives and cooperation mechanisms enabled by integrated digital platforms supporting Mobility-as-a-Service (MaaS). MaaS platforms streamline users’ abilities to plan, book, and pay for various mobility services through a unified digital interface and distribute those revenues between operators, thereby aligning physical infrastructure and digital coordination. As such, MHs form critical components of cyberphysical MaaS ecosystems, acting as strategic nodes for interactions among public entities, private mobility operators, and users.
@@
The deployment of MHs alongside MaaS schemes has been studied globally over the past decade. A notable example is the car-sharing program in Bremen, Germany (Karbaumer and Weltring,, 2025). Car-sharing stations known as “mobil.punkte” are established at easily accessible locations near major public transit stops, with supporting infrastructure such as bike racks and app-based booking systems. Sustainability-focused mobility hub designs have gained more attention with the maturation of electric vehicle (EV) technologies. The eHUBS project, supported by the European Regional Development Fund, investigated the deployment of intermodal hubs focusing on electric-powered modes in six pilot cities across North-West Europe (Amsterdam,, 2025). For a comprehensive review of mobility hub-focused pilot studies worldwide, readers can refer to Arnold et al., (2023).
@@
Designing systems incorporating MHs is no trivial task. Grigolon et al., (2025) reveals that travelers are willing to pay extra to use MHs when multiple modes are easily accessible through a well-designed physical location supported by a digital platform. Existing studies predominantly focus only on their role as physical transfer points (i.e. transit hubs), where the main objective is to optimally choose the hub locations that enhance accessibility (Petrović et al.,, 2019; Frank et al.,, 2021; Aydin et al.,, 2022). However, these studies often overlook the potential benefits of MHs on operator strategies. For example, MHs provide a location for stationing on-demand vehicles, fueling or charging electric vehicles, and provide a means for two operators to have controlled cost transfers (i.e. only subsidize trips that start or end at MHs). For example, an integration between Long Island Railroad (LIRR) and Uber may be costly if LIRR subsidized all Uber trips to all stations, but setting boundaries for specific stations through bundled tickets purchased for Uber and LIRR can produce partnerships in a more cost-effective manner that can suit both parties.
@@
Effective service integration requires joint operational planning among operators to improve traveler experiences. Recent studies, such as Xanthopoulos et al., (2024), partially address this gap by considering passenger preferences in determining optimal locations and capacities of mobility hubs. Nevertheless, these works still do not fully explore how such hub characteristics as location, capacity, pricing structures, and subsidies can directly shape operators’ key decisions, including fleet allocation, service area planning, and profitability. In reality, MHs can serve as powerful control levers within public cyberphysical platforms, critically influencing both demand-side traveler choices and supply-side operator strategies. This dual influence remains largely underexplored in the current literature.
@@
Recent advancements in MaaS modeling provide a feasible way in addressing the complex interactions between travelers and operators. Bilevel frameworks and many-to-many stable matching games (Liu and Chow,, 2024; Yao and Zhang,, 2024; Liu et al.,, 2024) have effectively captured market dynamics between travelers and operators. However, these models generally do not explicitly incorporate mobility hubs as integral network gateways. This gap highlights the need to systematically integrate mobility hub considerations into MaaS modeling frameworks, particularly in contexts involving complex interactions between public and private operators, such as providing first-mile and last-mile services to commuter rail stations. In addition, high computational complexity associated with the MaaS models prevents larger scale deployment, which is often associated with MH design scenarios.
@@
We fill this gap in the literature by proposing a mathematical model that optimizes the location and capacity of shared multimodal mobility hubs while maximizing travel utility in urban areas, accounting for multimodal trips and the cost transfers or subsidies between operators. In this context, operators participating in the MHs are part of a MaaS platform, which can be operated by a lead fixed route transit operator. To address computational complexity issues and scalability limitations inherent in existing multimodal flow assignment methods, we introduce the Perturbed Utility Route Choice (PURC)-based assignment game, based on the route choice framework from Fosgerau et al., (2022). Unlike path-enumeration approaches, the PURC framework uses link-level utilities to model multimodal route choices efficiently. This approach facilitates scalable and realistic MaaS market analyses involving mobility hubs. A case study for the neighborhoods along three LIRR stations in Suffolk county, NY, illustrates the application of the proposed mobility hub design model on a real-world scale. In addition, the case study provides valuable insights in how a MH can act as leverage in controlling the amount of subsidies in improving transit station usage.
@@ 2.1 MaaS platform models
MaaS platforms integrate various transportation modes into bundles that enable more seamless trips involving multiple service operators. A MaaS platform allows users to purchase trips consisting of combinations of mobility options in a single transaction. The pricing and service schemes have attracted considerable attention in recent years (van den Berg et al.,, 2022). Bertsimas et al., (2020) proposed a framework that jointly optimizes schedule frequencies and service pricing in a multi-modal transit network. Given a set of multimodal trip paths, service providers optimize their service frequencies to minimize system-wide wait time under budget constraints, while passengers select their preferred routes based on a nested logit model.
@@
Alternatively, two-sided market matching mechanism is adopted to model MaaS service design problems. Djavadian and Chow, (2017) pioneered the two-sided market mechanism into MaaS modeling. Rasulkhani and Chow, (2019) and Pantelidis et al., (2020) expanded the stable matching approach with transferable utilities to model the interaction of fixed-line service providers and travelers as a two-sided market, first as a many-to-one assignment game and then as a many-to-many assignment game. Travelers select their desired paths, which may involve multiple operators, while service operators determine whether to operate a service link and set service prices. Liu and Chow, (2024) further extended this framework to incorporate mobility-on-demand (MOD) services with congested links for accessing MOD. An exact solution algorithm was proposed to identify stable outcomes and determine the level of subsidy required from the platform to the operators to stabilize empty cores, highlighting the additional resources needed to operate MaaS platforms. Building on the many-to-many matching framework, Yao and Zhang, (2024) proposed a different design in which the MaaS platform acts as an intermediary which purchases capacity from service providers and offers service bundles to travelers with OD-based pricing. Additionally, non-MaaS travelers are considered within the congested network.
@@
Bilevel structures are also frequently used to formulate MaaS platform design problems. The upper level typically represents decisions made by the service platform or providers, while the lower level captures the decisions of other participants, such as travelers. Xi et al., 2024a proposed a single-leader multi-follower game (SLMFG) in which the MaaS platform is the leader, and travelers and service providers are followers. The platform determines service bundles and pricing, while travelers and service providers choose their participation to optimize their respective objectives. Building on this work, Xi et al., 2024b expanded the framework into a multi-leader multi-follower game (MLMFG) that considers competition among platforms in an electric MaaS (E-MaaS) ecosystem. Huang et al., (2024) used an SLMFG to model the interaction between a regulator and multiple service providers: the regulator provides path-based subsidies at the upper level, while non-cooperative service providers set link prices to maximize their own profits at the lower level. Pinto et al., (2020) adopted a bilevel structure to address a resource allocation problem, designing a multimodal system with autonomous vehicle (AV) fleets while capturing time-dependent mode choice behavior. Similarly, Bandiera et al., (2024) formulated a bilevel problem in which service operators determine optimal strategies at the upper level and traveler behaviors are modeled at the lower level. Instead of a centralized operator, each service provider maximizes its own profit, resulting in a Nash equilibrium at the upper level.
@@
The deterministic assignment game in (Liu and Chow,, 2024; Yao and Zhang,, 2024) was extended to a probabilistic approach by Liu et al., (2024) to reflect stochastic coalition choice. This choice is not identical to route choice. The assignment model adopts a similar path-based framework as the capacitated SUE model from Bell, (1995), but reflects the choices of the travelers and operators on which coalitions to form. This difference from a conventional route choice model is reflected in the added operator utility term in the lower-level objective function. Because the conventional SUE model requires path-based solutions, path set generation is a prerequisite of the solution process for the stochastic assignment game in Liu et al., (2024) as well. However, high quality path set generation is a non-trival task as discussed in (Prato,, 2009). Furthermore, the logit-based SUE in the lower level of a bilevel optimization makes it hard to integrate into a single level problem for solving to global optimality. As such, the path-based stochastic assignment game is not well-suited for large-scale applications.
@@ 2.2 Mobility Hubs
Mobility hubs can serve as key gateways that facilitate multi-modal mobility services. However, studies related to mobility hubs mainly focus on facility location and resource allocation, i.e. aspects of traditional transit hub design in which interactions between operators are ignored. Previous research has primarily focused on locating mobility hubs to improve accessibility and reduce social inequity, with the involved facilities typically centered around a single mode. Caggiani et al., 2020a and Caggiani et al., 2020b proposed models to optimally place bike-share stations to enhance spatial fairness in mobility access. Similarly, Duran-Rodas et al., (2021) developed a model to site bike-share stations with the goal of minimizing spatial inequity. Other studies on hub location decisions also treat social welfare as a central research focus (Banerjee et al.,, 2020; Aydin et al.,, 2022). Frank et al., (2021) improve rural accessibility by using mobility hubs as gateways for multimodal trip itineraries.
Although numerous models have been proposed for designing mobility hubs, only a few studies consider the interactions among mobility service participants. Nair and Miller-Hooks, (2014) was a pioneer in developing hub location models that incorporate traveler mode choices. They proposed a bilevel framework, with the hub operator as the upper-level decision maker and travelers as the lower-level decision makers. Ma et al., (2019) provide a ridesharing operation strategy that prioritize interconnections with transit networks. Though not specifically designed for mobility hub application, the case study uses LIRR stations as key points for deploying proposed ridesharing dispatch and fleet repositioning strategy, showing the potential of using multimodal services for last-mile trips. Xanthopoulos et al., (2024) also demonstrate the importance of considering user preferences in determining optimal hub locations and capacities. Their framework decomposes hub decisions and traveler choices into multiple modules, and a customized metaheuristic was developed to apply the framework at a city scale.
@@ 2.3 Research gaps and our contributions
Although numerous studies have developed models for designing multimodal service networks and mobility hubs, explicitly integrating mobility hubs into MaaS optimization frameworks remains a significant and under-explored area. Nair and Miller-Hooks, (2014) and Xanthopoulos et al., (2024) captured the interaction between operator resource allocation decisions and traveler path choices. However, such analyses primarily treat hubs as mere transfer points, without explicitly considering their strategic role in operator decision-making processes. Specifically, service pricing and subsidy between operators are not considered key decision variables, even though they play crucial roles in influencing traveler choices. Xi et al., 2024a ; Xi et al., 2024b proposed frameworks that leverage service capacity and pricing strategies in designing MaaS ecosystems. However, these frameworks do not consider the spatial aspects of resource allocation decisions and traveler itineraries, nor do they incorporate physical infrastructure when deciding service strategies.
The frameworks proposed by Liu and Chow, (2024) and Yao and Zhang, (2024) effectively capture the complex interactions among the MaaS platform, service providers, and travelers while considering flexible service strategies in multimodal networks. Nonetheless, the role of mobility hubs as key levers in MaaS ecosystem design remains unaddressed. Additionally, the above-mentioned studies often require substantial computational resources, hindering their deployments on a city-level scale. These gaps call for further modeling approaches capable of capturing strategic interactions among operators and travelers, enhanced by more scalable computational methods. A path-based stochastic assignment game approach (Liu et al.,, 2024; Liu,, 2024) addresses some of the scalability and capturing of heterogeneous preferences, but faces others with path generation requirements.
This study addresses these gaps by rigorously integrating MH characteristics into MaaS market models, thereby contributing to both theoretical advancement and practical implementation strategies. The main contributions of this study are summarized as follows:
@@
    We propose a bilevel mathematical model for the general MaaS platform design problem, given the network structure. The model aims to maximize MaaS platform revenue or social welfare while capturing the two-sided matching between mobility hub costs and traveler utility at the lower level. The upper-level decision variable is the service subsidy between operators through transfers at the MH, while the lower-level decision variables are link flow and mobility service node capacities.
@@
    We adopt a random utility model (RUM) based formulation to capture the choices of the coalitions between traveler and operator choices in the lower level as proposed by Liu, (2024). We extend it to a link-based formulation using the PURC approach proposed by Fosgerau et al., (2022). The approach offers improved scalability while still capturing link flows jointly determined by service operators and travelers, similar to other RUM formulations (e.g., stochastic user equilibrium, SUE).

### v2

@@ 1 Introduction
Emergent cyberphysical mobility hubs (MHs) are a viable response to these challenges. MHs are defined as ”multimodal transport nodes that facilitate intermodal transfers by providing different mobility options in close proximity” (Miramontes et al., 2017). While ”transit hubs” have existed for decades as a means to integrate land use and public transport as a type of ”transit-oriented development” (see Weustenenk and Mingardo (2023)), they differ from MHs in the literature in the last decade, which feature higher levels of both physical and digital integration of mobility services (Arias-Molinares et al., 2023), particularly services that feature real-time information. These MHs often involve multiple operators (or even non-mobility-related goods and service providers, such as fueling), necessitating distinct incentives and cooperation mechanisms enabled by integrated digital platforms supporting Mobility-as-a-Service (MaaS). MaaS platforms streamline users’ abilities to plan, book, and pay for various mobility services through a unified digital interface and distribute those revenues between operators, thereby aligning physical infrastructure and digital coordination. As such, MHs form critical components of cyberphysical MaaS ecosystems, acting as strategic nodes for interactions among public entities, private mobility operators, and users.
@@
The deployment of MHs alongside MaaS schemes has been studied globally over the past decade. A notable example is the car-sharing program in Bremen, Germany (Karbaumer and Weltring, 2025). Car-sharing stations known as “mobil.punkte” are established at easily accessible locations near major public transit stops, with supporting infrastructure such as bike racks and app-based booking systems. Sustainability-focused mobility hub designs have gained more attention with the maturation of electric vehicle (EV) technologies. The eHUBS project, supported by the European Regional Development Fund, investigated the deployment of intermodal hubs focusing on electric-powered modes in six pilot cities across North-West Europe (Amsterdam, 2025). For a comprehensive review of mobility hub-focused pilot studies worldwide, readers can refer to Arnold et al. (2023).
@@
Designing systems incorporating MHs is no trivial task. Grigolon et al. (2025) reveals that travelers are willing to pay extra to use MHs when multiple modes are easily accessible through a well-designed physical location supported by a digital platform. Existing studies predominantly focus only on their role as physical transfer points (i.e. transit hubs), where the main objective is to optimally choose the hub locations that enhance accessibility (Petrović et al., 2019; Frank et al., 2021; Aydin et al., 2022). However, these studies often overlook the potential benefits of MHs on operator strategies. For example, MHs provide a location for stationing on-demand vehicles, fueling or charging electric vehicles, and provide a means for two operators to have controlled cost transfers (i.e. only subsidize trips that start or end at MHs). For example, an integration between Long Island Railroad (LIRR) and Uber may be costly if LIRR subsidized all Uber trips to all stations, but setting boundaries for specific stations as designated MHs through bundled tickets purchased for Uber and LIRR can produce partnerships in a more cost-effective manner that can suit both parties. The problem becomes even more complex should other mobility provides wish to participate, i.e. how should LIRR consider a common platform linking with both Uber and Lyft, how should subsidies to these operators be designed, and which stations should be selected as MHs?
@@
Effective service integration requires joint operational planning among operators to improve traveler experiences. Recent studies, such as Xanthopoulos et al. (2024), partially address this gap by considering passenger preferences in determining optimal locations and capacities of mobility hubs. Nevertheless, these works still do not fully explore how such hub characteristics as location, capacity, pricing structures, and subsidies can directly shape operators’ key decisions, including fleet allocation, service area planning, and profitability. In reality, MHs can serve as powerful control levers within public cyberphysical platforms, critically influencing both demand-side traveler choices and supply-side operator strategies. This dual influence remains largely underexplored in the current literature.
@@
Recent advancements in MaaS modeling provide a feasible way in addressing the complex interactions between travelers and operators. Bilevel frameworks and many-to-many stable matching games (Liu and Chow, 2024; Yao and Zhang, 2024; Liu et al., 2025) have effectively captured market dynamics between travelers and operators. However, these models do not explicitly incorporate mobility hubs as integral network gateways. This gap highlights the need to systematically integrate mobility hub considerations into MaaS modeling frameworks, particularly in contexts involving complex interactions between multiple public and private operators, such as providing first-mile and last-mile services to commuter rail stations. In addition, high computational complexity associated with the MaaS models prevents larger scale deployment, which is often associated with MH design scenarios that must link both neighborhood subnetworks with larger regional subnetworks.
@@
We fill this gap in the literature by proposing a mathematical model that optimizes the capacity allocation of shared multimodal mobility hubs while maximizing travel utility, accounting for multimodal trips and the cost transfers and subsidies between operators. In this context, operators participating in the designated MHs are part of a MaaS platform, which can be operated by a lead fixed route transit operator. To address computational complexity issues and scalability limitations inherent in existing multimodal flow assignment methods, we introduce the Perturbed Utility Route Choice (PURC)-based assignment game, based on the route choice framework from Fosgerau et al. (2022). Unlike path-enumeration approaches, the PURC framework uses link-level utilities to model dispersed multimodal route choices efficiently. This approach facilitates scalable and realistic MaaS market analyses involving mobility hubs. A case study for the neighborhoods along three LIRR stations in Suffolk County, NY, illustrates the application of the proposed mobility hub design model on a real-world scale. In addition, the case study provides valuable insights to how a MH can act as leverage in controlling the amount of subsidies used to improve transit station usage.
@@ 2.1 MaaS platform models
MaaS platforms integrate various transportation modes into bundles that enable more seamless trips involving multiple service operators. A MaaS platform allows users to purchase trips consisting of combinations of mobility options in a single transaction. The pricing and service schemes have attracted considerable attention in recent years (van den Berg et al., 2022). Bertsimas et al. (2020) proposed a framework that jointly optimizes schedule frequencies and service pricing in a multi-modal transit network. Given a set of multimodal trip paths, service providers optimize their service frequencies to minimize system-wide wait time under budget constraints, while passengers select their preferred routes based on a nested logit model. Hörcher and Graham (2020) compared how different MaaS subscription policies affect system performance, specifically regarding net welfare, externalities, and operator objectives. The study involves three levels of decision making: car ownership, MaaS package selection, and daily mode choice. Each level is modeled using utility maximization principles with nonlinear externalities. They found that certain pricing policies can lead to unexpected negative outcomes by encouraging the overconsumption of services.
@@
Alternatively, two-sided market matching mechanism is adopted to model MaaS service design problems. Djavadian and Chow (2017) pioneered the two-sided market mechanism into MaaS modeling. Chen and Valant (2023) studied the convergence properties of these two-sided markets with a focus on ride-hail matching. Rasulkhani and Chow (2019) and Pantelidis et al. (2020) expanded the stable matching approach with transferable utilities to model the interaction of fixed-line service providers and travelers as a two-sided market, first as a many-to-one assignment game and then as a many-to-many assignment game. Travelers select their desired paths, which may involve multiple operators, while service operators determine whether to operate a service link and set service prices. Liu and Chow (2024) further extended this framework to incorporate mobility-on-demand (MOD) services with congested links (i.e. macroscopic matching impedance functions) for accessing MOD services. An exact solution algorithm was proposed to identify stable outcomes and determine the level of subsidy required from the platform to the operators to stabilize empty cores, highlighting the additional resources needed to operate MaaS platforms. Building on the many-to-many matching framework, Yao and Zhang (2024) proposed a different design in which the MaaS platform acts as an intermediary which purchases capacity from service providers and offers service bundles to travelers with OD-based pricing. Additionally, non-MaaS travelers are considered within the congested network.
@@
Bilevel structures are also frequently used to formulate MaaS platform design problems. The upper level typically represents decisions made by the service platform or providers, while the lower level captures the decisions of other participants, such as travelers. Xi et al. (2024a) proposed a single-leader multi-follower game (SLMFG) in which the MaaS platform is the leader, and travelers and service providers are followers. The platform determines service bundles and pricing, while travelers and service providers choose their participation to optimize their respective objectives. Building on this work, Xi et al. (2024b) expanded the framework into a multi-leader multi-follower game (MLMFG) that considers competition among platforms in an electric MaaS (E-MaaS) ecosystem. Huang et al. (2024) used an SLMFG to model the interaction between a regulator and multiple service providers: the regulator provides path-based subsidies at the upper level, while non-cooperative service providers set link prices to maximize their own profits at the lower level. Pinto et al. (2020) adopted a bilevel structure to address a resource allocation problem, designing a multimodal system with autonomous vehicle (AV) fleets while capturing time-dependent mode choice behavior. Similarly, Bandiera et al. (2024) formulated a bilevel problem in which service operators determine optimal strategies at the upper level and traveler behaviors are modeled at the lower level. Instead of a centralized operator, each service provider maximizes its own profit, resulting in a Nash equilibrium at the upper level.
@@
The deterministic assignment game in (Liu and Chow, 2024; Yao and Zhang, 2024) was extended to a probabilistic approach by Liu et al. (2025) to reflect stochastic coalition choice. The assignment model adopts a similar path-based route choice framework as the capacitated SUE model from Bell (1995), but the flows reflect the joint choices of the travelers and operators on which coalitions to match. This difference from a conventional route choice model is reflected in the added operator utility term in the lower-level objective function. Because the conventional SUE model requires path-based solutions, path set generation is also a prerequisite of the solution process for the stochastic assignment game in Liu et al. (2025). However, high quality path set generation is a non-trival task as discussed in (Prato, 2009). Furthermore, the logit-based SUE in the lower level of a bilevel optimization makes it hard to integrate into a single level problem for solving to global optimality. As such, the path-based stochastic assignment game is not well-suited for large-scale applications.
@@ 2.2 Mobility Hubs
Mobility hubs can serve as key gateways that facilitate multi-modal mobility services. However, studies related to mobility hubs mainly focus on facility location and resource allocation, i.e. aspects of traditional transit hub design in which interactions between operators are ignored. Previous research has primarily focused on locating mobility hubs to improve accessibility and reduce social inequity, with the involved facilities typically centered around a single mode. Caggiani et al. (2020a, b) proposed models to optimally place bike-share stations to enhance spatial fairness in mobility access. Similarly, Duran-Rodas et al. (2021) developed a model to site bike-share stations with the goal of minimizing spatial inequity. Other studies on hub location decisions also treat social welfare as a central research focus (Banerjee et al., 2020; Aydin et al., 2022). Frank et al. (2021) improve rural accessibility by using mobility hubs as gateways for multimodal trip itineraries. Ren and Chow (2025) integrate synthetic data with observed data to evaluate the impacts of a MH on traveler disutilities.
Although numerous models have been proposed for designing mobility hubs, only a few studies consider the interactions among mobility service participants. Nair and Miller-Hooks (2014) was a pioneer in developing hub location models that incorporate traveler mode choices. They proposed a bilevel framework, with the hub operator as the upper-level decision maker and travelers as the lower-level decision makers. Ma et al. (2019) provide a ridesharing operation strategy that prioritize interconnections with transit networks. Though not specifically designed for mobility hub application, the case study uses LIRR stations as key points for deploying proposed ridepooling dispatch and fleet repositioning strategy. The results show that the presence of a trunk commuter transit network would redistribute optimal ridepooling services to the last mile to increase effective service capacity by mitigating the ”Wild Goose Chase” phenomenon (see also Ouyang and Yang (2023)). Xanthopoulos et al. (2024) also demonstrate the importance of considering user preferences in determining optimal hub locations and capacities. Their framework decomposes hub decisions and traveler choices into multiple modules, and a customized metaheuristic was developed to apply the framework at a city scale.
@@ 2.3 Research gaps and our contributions
Although numerous studies have developed models for designing multimodal service networks and mobility hubs, explicitly integrating mobility hubs into MaaS optimization frameworks remains a significant and under-explored area. Nair and Miller-Hooks (2014) and Xanthopoulos et al. (2024) captured the interaction between operator resource allocation decisions and traveler path choices. However, such analyses primarily treat hubs as mere transfer points, without explicitly considering their strategic role in operator decision-making processes. Specifically, service pricing and subsidy between operators are not considered key decision variables, even though they play crucial roles in influencing traveler choices. Xi et al. (2024a, b) proposed frameworks that leverage service capacity and pricing strategies in designing MaaS ecosystems. However, these frameworks do not consider the spatial aspects of resource allocation decisions and traveler itineraries, nor do they incorporate physical infrastructure when deciding service strategies.
The frameworks proposed by Liu and Chow (2024) and Yao and Zhang (2024) effectively capture the complex interactions among the MaaS platform, service providers, and travelers while considering flexible service strategies in multimodal networks. Nonetheless, the role of mobility hubs as key levers in MaaS ecosystem design remains unaddressed. Additionally, the above-mentioned studies often require substantial computational resources, hindering their deployments on a city-level scale. These gaps call for further modeling approaches capable of capturing strategic interactions among operators and travelers, enhanced by more scalable computational methods. A path-based stochastic assignment game approach (Liu et al., 2025) addresses some of the scalability and capturing of heterogeneous preferences, but faces others with path generation requirements.
This study addresses these gaps by rigorously integrating MH characteristics as platform control levers into MaaS market models, thereby contributing to both theoretical advancement and practical implementation strategies. The main contributions of this study are summarized as follows:
@@
    We propose a bilevel mathematical model for the general MaaS platform design problem, given the network structure. The model aims to maximize MaaS platform revenue (if private) or social welfare (if public) while capturing the two-sided matching between mobility hub costs and traveler utility at the lower level. The upper-level decision variables are the OD-based traveler-facing pricing that grants access to the MaaS services and operator-dedicated subsidies, while the lower-level decision variables are link flows and mobility service node capacities. The proposed model involves heterogeneous operator participation and the competition mechanism, using subsidies as a lever to transfer utility between travelers and operators with heterogeneous service elements.
@@
    We adopt a random utility model (RUM) based formulation to capture the choices of the coalitions between traveler and operator choices in the lower level as proposed by Liu et al. (2025). We extend it to a link-based formulation using the PURC approach proposed by Fosgerau et al. (2022). The approach offers improved scalability while still capturing link flows jointly determined by service operators and travelers, similar to other RUM formulations (e.g., stochastic user equilibrium, SUE).